what is the equation y=-.5x-4 in slope-intercept form for a line perpendicular to the graph of the equation that passes through the x-intercept of that line?
Answer by TimothyLamb(4379)

L1:
y = -.5x - 4
same as:
y = (-1/2)x - 4
slope (m) = (-1/2)
x-intercept (a) = 4(-2) = (-8, 0)
---
L2:
perpendicular lines have negative reciprocal slopes:
mp = 2
use the point-slope form with x-intercept of L1:
y - y1 = mp(x - x1)
y = 2(x + 8)
Answer:
eqn. of perp. line in slope-intercept form:
y = 2x + 16
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php